OpenStudy (anikate):

Someone pl help me with sq3(sq15+sq18)= sq=square root

OpenStudy (anonymous):

\[\huge \sqrt[3]{\sqrt{15}+\sqrt{18}}\] This your question?

OpenStudy (anikate):

no just a minute I will type that way

OpenStudy (anikate):

\[_{0}^{3}(_{15}^{?}+_{18}^{?}\]

OpenStudy (anikate):

sry uh I dont get it how to put it

OpenStudy (anikate):

wait let me try in different way

OpenStudy (anonymous):

you can use the draw feature too

OpenStudy (anonymous):



OpenStudy (anikate):

sq3